OpenStudy (anonymous):

What is the solution set for |2x – 3| = 13

OpenStudy (anonymous):

x=8 and -5

jimthompson5910 (jim_thompson5910):

|2x – 3| = 13 2x – 3 = 13 or 2x – 3 = -13 2x = 13+3 or 2x = -13+3 2x = 16 or 2x = -10 x = 16/2 or x = -10/2 x = 8 or x = -5